# ULPT: Prompt Tuning with Ultra-Low-Dimensional Optimization

### v1

@@ 5 Conclusion
In this paper, we propose Ultra-Low-Dimensional Prompt Tuning (ULPT), a novel parameter-efficient prompt tuning method that achieves superior performance across diverse NLP tasks with significantly fewer trainable parameters. ULPT decouples prompt embeddings from the model’s dimensionality, optimizing in a low-dimensional space and projecting into the model’s embedding space by a frozen random projection.
Our research offers future opportunities for large-scale LLM customizations, as efficient storage of task-specific models is increasingly critical.
Appendix A Appendix
Appendix B Dataset Details
Table 5: Dataset information and statistics.
@@ Appendix B Dataset Details
We present detailed information for the 21 NLP tasks in Table 5. Following previous work (Wang et al., 2023; Shi and Lipani, 2024), we preprocess the labels for classification and multiple-choice tasks into a single-token label (e.g., 0, 1, 2, …) to simplify evaluation. For MRQA, the model generates an answer containing a sequence of tokens.
@@
Based on the training set size, the tasks can be roughly categorized into three scales: small ($<$10K samples), medium (10–100K samples), and large ($>$100K samples). Notably, SuperGLUE contains small training sets, and is generally considered more challenging than GLUE, making it more susceptible to overfitting due to its limited samples. By contrast, MRQA and the tasks in the “Others” category consist of more complex tasks, likely requiring more parameters to capture their difficulty.
Appendix C Theoretical Results
@@ C.1 Proof of Theorem 1
These are the common assumptions used to show the optimization process in deep learning . In addition, we also impose an assumptions on the projection matrix and the scaling vector ${\bm{s}}$.
@@
Let ^$\hat{\mathcal{L}}(\hat{\bm{x}})$ be the loss function with our ULPT approach, where $\hat{\bm{x}}\in\mathbb{R}^{nr+2d}$ is the concatenation of shift/scale embeddings and the ultra-low-dimensional prompt embeddings. ^$\hat{\mathcal{L}}(\hat{\bm{x}})$ is $\beta^{\prime}$-Lipschitz w.r.t. $\hat{\bm{x}}$ for some $\beta^{\prime}>0$.
@@
|^$\left|\nabla\hat{\mathcal{L}}(b_{k}^{(1)})-\nabla\hat{\mathcal{L}}(b_{k}^{(2)})\right|=$ $\left|\sum_{i=1}^{n}\left(\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(1)})-\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(2)})\right)\right|$ (19)
@@
|^$\left|\nabla\hat{\mathcal{L}}(s_{k}^{(1)})-\nabla\hat{\mathcal{L}}(s_{k}^{(2)})\right|=$ $\left|\sum_{i}\left({\bm{z}}_{i}^{\top}{\bm{P}}_{:,k}\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(1)})-{\bm{z}}_{i}^{\top}{\bm{P}}_{:,k}\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(2)})\right)\right|$ (23)
@@
|^$\left|\nabla\hat{\mathcal{L}}(z_{i,k}^{(1)})-\nabla\hat{\mathcal{L}}(z_{i,k}^{(2)})\right|=$ $\left|{\bm{P}}_{k,:}\operatorname{diag}({\bm{s}})\frac{\partial{\mathcal{L}}}{\partial\hat{\bm{e}}_{i}^{(1)}}-{\bm{P}}_{k,:}\operatorname{diag}({\bm{s}})\frac{\partial{\mathcal{L}}}{\partial\hat{\bm{e}}_{i}^{(2)}}\right|$ (30)

### v2

@@ 5 Conclusion
We introduce ULPT, a novel parameter-efficient method that learns task-specific prompts in a low-dimensional space and projects them into the model space using a frozen random matrix with learned shift and scale vectors. ULPT achieves competitive or superior performance compared with state-of-the-art parameter-efficient fine-tuning methods while using dramatically fewer trainable parameters.
6 Limitations
We have comprehensively evaluated ULPT on both encoder–decoder T5 models and decoder-only Bloomz and Llama models on various NLP and reasoning datasets, demonstrating its parameter efficiency and competitive performance. However, due to the limitation in computational resource, we have not explored its effectiveness on very large language models (ranging from tens to hundreds of billions of parameters). We anticipate that, for such large LLMs, parameter-efficient methods like ULPT are particularly suitable for lightweight customization, such as adapting generation style or output formatting Liu et al. (2024b), rather than unlocking new capabilities which are better addressed in pre-training. This highlights the advantage of ULPT in scenarios that require ultra-low parameter usage without compromising effectiveness.
Acknowledgments
We thank the reviewers and area chairs for their efforts. The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Amii Fellow Program, the Canada CIFAR AI Chair Program, a donation from DeepMind, and the Digital Research Alliance of Canada (alliancecan.ca).
Appendix A Experimental Details
A.1 Datasets Statistics for Language Understanding Tasks
@@
Table 4: Dataset information and statistics for the main experiments in Section 4.1.
@@
We present detailed information for the 21 natural language understanding tasks in Table 4. Following previous work (Wang et al., 2023; Shi and Lipani, 2024), we preprocess the labels for classification and multiple-choice tasks into a single-token label (e.g., 0, 1, 2, …) to simplify evaluation. For MRQA, the model generates an answer containing a sequence of tokens.
@@
Based on the training set size, the tasks can be roughly categorized into three scales: small (<<10K samples), medium (10–100K samples), and large (>>100K samples). Notably, SuperGLUE contains small training sets, and is generally considered more challenging than GLUE, making it more susceptible to overfitting due to its limited samples. By contrast, MRQA and the tasks in the “Others” category consist of more complex tasks, likely requiring more parameters to capture their difficulty.
A.2 Training Details for Language Understanding Tasks
In our experiment for $21$ language understanding tasks (Section 4.1), we use a batch size of $16$ and a default learning rate of 6e16e{-1} with AdamW. The learning rate follows a linear schedule, warming up for $500$ steps and then decaying linearly to 0. We set a maximum sequence length of $256$ for most tasks, except for SuperGLUE-MultiRC being $348$ and MRQA being $512$. ULPT is trained on all tasks for up to $100,000$ steps. Performance is evaluated every $1,000$ steps, with the best checkpoint selected based on the validation set.
A.3 Implementation and Training Details for Reasoning Tasks
In Section 4.2, we evaluate ULPT on Llama 3.2 Meta (2024), a decoder-only model in 1B ($d=2048$) and 3B ($d=3072$) scales. We consider two reasoning tasks, GSM8K Cobbe et al. (2021) for math reasoning and MBPP Austin et al. (2021) for code generation.
As baselines, we include in-context learning (4-shot demostrations sampled from each training set), LoRA and VeRA adapters with ranks $r\in\{1,4,8\}$, FourierFT with three frequency configurations $n\in\{128,512,1024\}$. Vanilla prompt tuning and ULPT both run with $10$ learnable tokens, since extending the length yields no additional benefit on Llama. We evaluate ULPT on three rank configurations of 2, 64 and 256. All methods share a batch size of 4, with training for 3 epochs on GSM8K and 10 epochs on MBPP. We set learning rates to 1e-3 for LoRA and VeRA, and 1e-2 for ULPT, prompt tuning and FourierFT. At inference, we apply greedy decoding with the maximum generation length of $1024$ tokens.
MRQA Others
Method #Param NQ HQA SQA NewsQA Avg. WG Yelp SciTail PAWS Avg.
Fine-tuning 220M 75.1 77.5 81.1 65.2 74.7 61.9 96.7 95.8 94.1 87.1
Adapter 1.9M 74.2 77.6 81.4 65.6 74.7 59.2 96.9 94.5 94.3 86.2
BitFit 280K 70.7 75.5 77.7 64.1 72.0 57.2 94.7 94.7 92.0 84.7
LoRA 3.8M 72.4 62.3 72.5 56.9 66.0 58.2 97.1 94.7 94.0 86.0
SPoT 76.8K 68.2 74.8 75.3 58.2 69.1 50.4 95.4 91.2 91.1 82.0
ATTEMPT 232K 70.4 75.2 78.5 62.8 71.4 57.6 96.7 93.1 92.1 84.9
PT 76.8K 70.0 74.7 75.3 63.0 70.8 49.6 95.6 92.0 57.9 73.8
DPT(r=10) 9.0K 71.3 75.5 76.3 63.5 71.7 49.6 96.1 95.6 92.2 83.4
DPT (r=256)  [^2] K 71.4 76.0 77.6 64.2 72.3 49.6 96.3 95.2 55.8 74.2
ULPT (r=2) 1.7K 67.20.2 74.00.1 71.70.2 61.40.1 68.6 49.50.2 95.60.1 93.00.9 90.40.2 82.1
Appendix B Additional Results
B.1 Distribution of Pretrained Embeddings
Figure 7: Distribution of randomly selected dimensions from the pretrained token embeddings of T5-base.
Following our pilot study in Section 3.2, we present the distribution of the same randomly selected dimensions from the pretrained token embeddings of the T5-base model. We observe that the variation in these pretrained embeddings is smaller compared with the learned soft embeddings from the CoLA and SST-2 tasks.
B.2 Additional Results on T5 Model
We evaluate ULPT on the MRQA dataset and four additional tasks in the “Others” category. Following the standard practice on these benchmarks Wang et al. (2023); Shi and Lipani (2024), we run ULPT three times with different seeds and report the mean and standard deviation. This setup also allows us to assess ULPT’s sensitivity to the random initialization of the up-projection matrix.
The results are shown in Table 5. Unlike GLUE and SuperGLUE performance, ULPT exhibits consistent improvement when the rank is higher. This is probably because these tasks are more challenging for the T5-base model, which aligns with the observation that full-model fine-tuning outperforms parameter-efficient methods on these tasks. Nevertheless, ULPT achieves competitive performance (slightly worse than the best-performing DePT approach), while saving parameters by multiple folds.
B.3 Additional Analysis on Bloomz Family
We further test our hypothesis from Section 4.3: under a fixed parameter budget, longer low-dimensional prompts offer greater expressivity. To this end, we extend ULPT evaluation to Bloomz Muennighoff et al. (2023), a decoder-only model available in 560M, 1.7B, and 3B sizes, having hidden dimensions of 1024, 2048, and 2560 respectively.
For evaluation diversity, we select three mid-sized tasks from each task group: SST-2, HotpotQA, and Winogrande, providing assessment across classification, multi-hop reasoning, and commonsense reasoning. Given that Bloomz models are larger than the T5 series, we train up to 30K steps with a batch size of $4$, while keeping other hyperparamters the same as our main experiment.
We consider comparing ULPT with prompt tuning under different parameter budgets for text generation. Specifically, we vary the rank of ULPT from $2$ to $256$ while fixing the length $n=100$. For full-dimensional prompt tuning, the token number is adjusted to match the parameter count, resulting in fewer virtual tokens than ULPT.
The results in Table 6 show that ULPT consistently outperforms full-dimensional prompt tuning across all model sizes and tasks given a fixed parameter budget. These results support our earlier analysis (Section 4.3) that allocating the parameter budget toward longer sequence lengths rather than higher embedding dimensions yields greater expressivity.
B.4 Inference Overhead Analysis
Runtime Setting Llama 1B Llama 3B
Vanilla PT embeddings loading $0.64\pm 0.04$ $0.91\pm 0.04$
ULPT up-projection ($r=2$) $0.56\pm 0.06$ $0.59\pm 0.04$
ULPT up-projection ($r=64$) $1.43\pm 0.09$ $1.87\pm 0.06$
ULPT up-projection ($r=256$) $4.09\pm 0.10$ $5.80\pm 0.16$
Decoding $1481.15\pm 64.26$ $2536.67\pm 42.14$
Table 7: Runtime (ms) for ULPT prompt reconstruction.
Model No adaptation ULPT
Llama 1B $82.76\pm 0.33$ $82.71\pm 0.33$
Llama 3B $48.74\pm 0.25$ $48.70\pm 0.22$
Table 8: Decoding speed (tokens/second)
ULPT’s ideal use case is for massive LLM customization, where adapters are frequently loaded to a model. Unlike vanilla prompt tuning, ULPT reconstructs prompt embeddings from a low-dimensional space at inference time. We empirically compare its up-projection runtime against vanilla PT over 100 runs, shown in Table 7. We find the overhead is negligible relative to the decoding time on GSM8K .
ULPT follows the paradigm of prompt tuning to prepend tokens. We also measure end-to-end decoding speed using ULPT (rank 2 and 100 prompt tokens) on a single NVIDIA A6000 48G GPU with the Huggingface generation API. As seen in Table 8, the decoding speed is similar with or without the additional prompt tokens.
Appendix C Theoretical Analysis
@@ C.1 Proof of Theorem 1
These are the common assumptions used to show the optimization process in deep learning (Karimi et al., 2016; Mei et al., 2020). In addition, we also impose an assumptions on the projection matrix and the scaling vector $\bm{s}$}.
@@
[…]
@@
$\left|\nabla\hat{\mathcal{L}}(b_{k}^{(1)})-\nabla\hat{\mathcal{L}}(b_{k}^{(2)})\right|=$ $\left|\sum_{i=1}^{n}\left(\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(1)})-\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(2)})\right)\right|$ (19)
@@
$\left|\nabla\hat{\mathcal{L}}(s_{k}^{(1)})-\nabla\hat{\mathcal{L}}(s_{k}^{(2)})\right|=$ $\left|\sum_{i}\left({\bm{z}}_{i}^{\top}{\bm{P}}_{:,k}\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(1)})-{\bm{z}}_{i}^{\top}{\bm{P}}_{:,k}\nabla{\mathcal{L}}(\hat{e}_{i,k}^{(2)})\right)\right|$ (23)
@@
$\left|\nabla\hat{\mathcal{L}}(z_{i,k}^{(1)})-\nabla\hat{\mathcal{L}}(z_{i,k}^{(2)})\right|=$ $\left|{\bm{P}}_{k,:}\operatorname{diag}({\bm{s}})\frac{\partial{\mathcal{L}}}{\partial\hat{\bm{e}}_{i}^{(1)}}-{\bm{P}}_{k,:}\operatorname{diag}({\bm{s}})\frac{\partial{\mathcal{L}}}{\partial\hat{\bm{e}}_{i}^{(2)}}\right|$ (30)